OpenStudy (anonymous):

I could use some help understanding thiss one please: Solve the linear system of equations 3x-2y=7 9x-6y=14 The choices are: A) 6,-4 B) 8,5 C) t,8t+5 D) No Solution I know it isn't A or B, but before I choose D, how do I do this with Choice C?

OpenStudy (anonymous):

PS- A and B are "Uniqe solutions" C is "Infinitely many solutions" if that means anything...

OpenStudy (anonymous):

Can you put them both into slope intercept form?

OpenStudy (anonymous):

I have no idea

OpenStudy (anonymous):

Please be aware, I suck at anything Math...lol

OpenStudy (anonymous):

Slope intercept form is the form y=mx+b, putting both of them in that form we have First one: 3x-2y=7 -2y=-3x+7 y=3/2x-7/2 second one: 9x-6y=14 -6y=-9x+14 y=3/2x-7/3 you can now see the lines have the same slope, and are therefore parallel which means they will never intersect. There is no solution

OpenStudy (anonymous):

Ok, I remember that formula now...lol However, how did you get y=3/2x-7/2?

OpenStudy (anonymous):

sorry, I just don't see how "C"" is in there...

OpenStudy (anonymous):

????